503what are the oxidation numbers of o and s
Answer #1

Oxidation number of O = -2

lets the oxidation number of S be x

use:

3* oxidation number (O) + 1* oxidation number (S) = net charge

3*(-2)+1* x = -2

-6 + 1 * x = -2

x = 4

So oxidation number of S = +4

Oxidation number of each element in SO3-2 is

O=-2

S=+4